OpenStudy (anonymous):

5(x – 1) + 3x = 7(x + 1)

OpenStudy (anonymous):

what is the soloution

OpenStudy (anonymous):

????

OpenStudy (stamp):

5(x - 1) + 3x = 7(x + 1) (5x - 5) + 3x = 7x + 7 8x - 5 = 7x + 7 You should be able to solve the rest and find a solution.

OpenStudy (anonymous):

i caint

OpenStudy (stamp):

Someone else will be better off helping you, I am sorry.

OpenStudy (anonymous):

ok thanks